A) you are given the following exchange rates
SF in Zurich: 1:6000 - 35SF/1$
£ in NY: 1:9810 - 50$/1 £

I) As a banker in London you would like to quote a £ cross-rate for the SF (i.e, £/1SF). If you anticipate to be the sole banker in the world quoting this cross-rate, what is your quote going to be? Explain
1
Expert's answer
2021-10-14T11:37:29-0400

Solution:

The banker will quote a cross rate of 0.9959.


This is calculated below:

The cross-rate formula is given by = "\frac{C}{B} = \frac{A}{B} \times \frac{B}{C}"


"\frac{£}{1SF} = \frac{£}{\$}\times \frac{\$}{SF}"


"=(\frac{1}{(\$\div£)ask }\times \frac{1}{(SF\div\$)ask}) \div (\frac{1}{(\$\div£)bid }\times \frac{1}{(SF\div\$)bid})"


"=(\frac{1}{1.9850}\times \frac{1}{1.6035}) \div (\frac{1}{1.9810 }\times \frac{1}{1.6000})"


"=\frac{0.3142}{0.3155} = 0.9959"


£/1SF = 0.9959